OpenStudy (anonymous):

15.) "where defined, find cos(theta) - sin(theta). simplify your answer."

OpenStudy (anonymous):

\[\cos(\Theta) - \sin(\Theta)\]

OpenStudy (anonymous):

that is one fat \[\theta\]

OpenStudy (anonymous):

what does it mean "find \[\sin(\theta)-\cos(\theta)\]?

OpenStudy (anonymous):

you want some other expression for this?

OpenStudy (anonymous):

i hav no clue

OpenStudy (anonymous):

i suppose so? that's all the instruction i got.

OpenStudy (anonymous):

oh oh oh i think i have seen these before. hold on

OpenStudy (anonymous):

you are supposed to write as a single trig function. i did one of these recently

OpenStudy (anonymous):

can you tell me the name of the topic?

OpenStudy (anonymous):

i had a nice recourse for these that made it easy. but i cannot find it. if you just want the answer it is here http://www.wolframalpha.com/input/?i=sin(x)+-+cos(x)

OpenStudy (anonymous):

but i had a nice explanation and method for getting these. i just cannot find it

OpenStudy (anonymous):

trigonometry?

OpenStudy (anonymous):

yeah i got that part. are you using a book? what is the section called?

OpenStudy (anonymous):

It's a placement practice test :)

OpenStudy (anonymous):

I got it off of some community college website.

OpenStudy (anonymous):

the question isnt clear to me.. when it says 'where defined' .. .. .. i dont see why is cant be straightforward and ask for a derivative or something

OpenStudy (anonymous):

what it is asking for is to write this as a single trig function. if it actually does not say that then it is a dumb question. did you see the answer i sent?

OpenStudy (anonymous):

it is of course always defined

OpenStudy (anonymous):

on the answer sheet. it says the answer is \[\cot(\theta) \cos(\theta)\]

OpenStudy (anonymous):

no. ALL IT SAYS is "where defined, find cos(theta)-sin(theta). simplify your answer"

OpenStudy (anonymous):

what???

OpenStudy (anonymous):

-_- that's why im confused.

OpenStudy (anonymous):

no no no that is not right.

OpenStudy (anonymous):

cause .. for me... they do not equal each other

OpenStudy (anonymous):

no they do not so forget about it . one answer is \[\sqrt{2}\sin(\theta-\frac{\pi}{4})\]

OpenStudy (anonymous):

i am trying to find the nice paper i had that showed how to do this, but i cannot. the answer you wrote is just straight up wrong so ignore it

OpenStudy (anonymous):

LoL :)

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

I guess there are a few things u might do if u want to get a different expression. U could use Sin/Cos = Tan Or square it -< sin^2 + cos^2 -2sin cos -> 1- 2 sin cos